write the equation is standard form for the line that contains the given point with the given slope.
(1,0); m=3

Answer by checkley71(8403)

Y=mX+b OR Y=3X+b OR SETTING Y=0 & X=1 WE GET 0=3*1+b OR b=-3
SO THE EQUATION IS Y=3X-3